OpenStudy (anonymous):

\(\ \Huge 36=\log_{2}x, solve for x? \)

OpenStudy (lgbasallote):

do you know how to change that into exponential form?

OpenStudy (lgbasallote):

@Study23 ...yes or no?

OpenStudy (anonymous):

2 to the what power equals 36?

OpenStudy (anonymous):

\(\ \Huge 2^{x} =36, \mathsf{how do I find x?} \)

hartnn (hartnn):

if \(\huge log_ab=c\) then \(\huge a^c=b\). u see where have u done the mistake ?

OpenStudy (lgbasallote):

you changed it wrong

OpenStudy (lgbasallote):

\[\huge \log_{base} (power) = exponent\] so for example you have \[\huge \log_2 4 = 2\] this translates to \[\huge \log 2^2 = 4\] got it?

OpenStudy (anonymous):

So for this problem, it would be \(\ \Huge x = 2^{36} ? \)

hartnn (hartnn):

yup, thats correct :)

OpenStudy (anonymous):

So some really big number?

hartnn (hartnn):

u wanted the exact value? then yes, very big number....else leave it in this form.

OpenStudy (lgbasallote):

you can do it by hand too

hero (hero):

Actually, x is not a really big number.

hero (hero):

5 < x < 6

OpenStudy (anonymous):

@hero ?

hero (hero):

\(2^5 = 32\) and \(2^6 = 64\) therefore, \(5 < x< 6\)

hartnn (hartnn):

hmmm....but,i got 2^36 as 68719476736

hero (hero):

I thought the original problem was \(2^x = 36\)

OpenStudy (anonymous):

Oh, no it wasn't...

hero (hero):

But you posted something to that effect earlier.